Solve the following LPPs using the simplex method




Max.Z=20x1 +10x2




Subject to:




5x1+4x2 < 250




2x1+5x2 < 150




x1, x2 > 0





1
Expert's answer
2022-04-05T01:51:47-0400

"\\left\\{ \\begin{array}{c}\t5x_1+4x_2+x_3=250\\\\\t2x_1+5x_2+x_4=150\\\\\\end{array} \\right. \\\\Z=20x_1+10x_2\\rightarrow \\max \\\\Simplex\\,\\,table:"



"The\,\,solution:\\x_1=50,x_2=0"